Week of 160502

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Maria (chair and minutes), Michal (ATLAS), Maarten (ALICE), Marcelo (LHCb), Lorena (Computing), Charles (DB), Marian (Monitoring and Networking), Maria (GGUS)
  • remote: Tommaso (CMS), Salvatore (CNAF), Dave (FNAL), Rolf (IN2P3), Sang Un (KISTI), Dimitri (KIT), Jens (NDGF), Dennis (NL-T1), Kyle (OSG), Di (TRIUMF)

Experiments round table:

  • ATLAS reports (raw view)
    • Activities:
      • MC15c, derivation production, and Heavy Ion reprocessing ongoing
      • re-validation of T0 spillover mechanism
    • Problems:
      • T1 DATADISKs getting full - BNL, SARA, INFN, FZK - data being rebalanced
      • BNL-SARA transfers are slow (GGUS:120957)
        • both BNL-CERN and BNL-SARA transfer slowness seems to be fixed
        • to be tested by DDM ops
      • transfers from FZK-LCG2_MCTAPE are failing with "Transfer canceled because the gsiftp performance marker timeout" (GGUS:121163)

  • CMS reports (raw view) -
    • Data taking with beam stopped due to the weasel (or whatever it was), switched to a CRAFT (Cosmic Run at 4 Tesla) on Saturday, expecting ~4 data at least.
    • we were really unfortunate, since just before the problem CMS had ramped the magnet back to full field, and we were eager to take data .... uff
    • that apart: production in full swing, no major problem
    • since yesterday a drop in the # of jobs, investigating ... (could be an effect of the end of batch#1 of requests)
    • for what concerns services, it was a very calm week.
    • A few sites (T2_PL_Swierk, T2_BE_IIHE, T2_UK_SouthGrid, T2_UK_SGrid_RALPP) have Phedex agents down
    • some files T1_ES_PIC -> reported as not existing at source (GGUS:121183).

  • ALICE -
    • High to very high activity
    • CASTOR incident starting Thu late evening
      • DAQ could no longer write to CASTOR
      • online alerted the CC operator at 02:56
      • CASTOR expert cured the problem by 04:00, thanks!
      • issue is understood and the devs are working on a fix

  • LHCb reports (raw view) -
    • Activity
      • Pre-staging of 2012 data for incremental stripping launched
    • Site Issues
      • T1:
        • CNAF - Files slowly staging causing a huge backlog. Under investigation.
        • GRIDKA - staging failing during the weekend . Under investigation

Sites / Services round table:

  • ASGC: NA
  • BNL: Eric couldn't join the meeting today
  • CNAF: NTR
  • FNAL: dCache upgraded to 2.13.
  • GridPP: NA
  • IN2P3: Intermittent gLExec error in CMS pilot jobs under investigation (GGUS:121026).
  • JINR: NA
  • KISTI: NTR
  • KIT:
    • On Tue early morning, we had lost both controllers of a disk storage cluster, which was used for several infrastructure services, one of which was to provide a mount to the Creams, that holds the gridmapdirs (GGUS:121044). Also, part of xrootd for Alice and the entire dgridsrm-fzk.gridka.de were down, too.
    • Last week's downtime for tape unfortunately had to be canceled due to before mentioned incident. We postponed it to next week Tue and Wed (whole days).
  • NDGF:
    • OPN Network intervention happening this evening. Short interruption of open connections (GOCDB)
    • dCache upgrade to 2.15.5 scheduled tomorrow (GOCDB)
  • NL-T1: NTR
  • NRC-KI: NA
  • OSG: NTR
  • PIC: NTR
  • RAL: NA
  • TRIUMF: NTR

  • CERN computing services:
    • Alarm ticket on Thursday for ATLAS for all CERN CREAM CEs.. turned out to be a known problem between Panda and CREAM (fixed by restarting Condor on APF).
  • CERN storage services: NA
  • CERN databases: NTR
  • GGUS: NTR
  • MW Officer: NA
  • Security: NA
  • Monitoring:
    • Draft reports for the April availability sent to the wlcg office
    • Ongoing issue with FAX US data missing from the XRootD dashboard
    • ATLAS has now understood GFAL issues so that ETF can go to production. Waiting for their green light.
  • Networking:
    • BNL-SARA link is now OK.
    • ASGC connectivity issues are progressing. There is better throughput, however, there is limited buffer which may be the cause of the issues. Investigations are still ongoing. If this is finally the cause, it will require an upgrade on the HW.

AOB: Reminder by MariaD: Tier1s were invited at the MWR and the the Ops Coord meetings on 16 & 17/3 to tell the e-group wlcg-ops-coord-wg-middleware at cern.ch whether they agree to install the pakiti client on their production service nodes, so that the versions of MW run at the site be known to authorised DNs of site managers taken from GOCDB and expert operations' supporters. Site replies received so far (CERN, JINR, Grif, RAL, NL_T1) on their intention to expand the use of pakiti client can be found here. All the input received is concatenated and attached to the next MW Readiness WG agenda of May 18th. Then we'll conclude on the issues.

Edit | Attach | Watch | Print version | History: r14 < r13 < r12 < r11 < r10 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r14 - 2016-05-02 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback